# The purpose of an estimator is what it does: Misspecification, estimands, and over-identification

### v4

@@ Abstract.
In over-identified models, misspecification—the norm rather than ex-
ception—fundamentally changes what estimators estimate. Different estimators imply different estimands rather than different efficiency for the same target. A review of recent applications of generalized method of moments in the American Economic Review suggests widespread acceptance of this fact: There is little formal specification testing and widespread use of estimators that would be inefficient were the model correct, including the use of “hand-selected” moments and weighting matrices. Motivated by these observations, we review and synthesize recent results on estimation under model misspecification, providing guidelines for transparent and robust empirical research. We also provide a new theoretical result, showing that Hansen’s J-statistic measures, asymptotically, the range of estimates achievable at a given standard error. Given the widespread use of inefficient estimators and the resulting researcher degrees of freedom, we thus particularly recommend the broader reporting of J-statistics.
@@ 1. Introduction
Scientific and policy questions in economics frequently hinge on quantities such as counterfactual outcomes or effects on welfare. Recovering them often requires strong modeling assumptions. At the same time, many economists concede that economic models are at best approximately true (White, 1982; Hansen and Sargent, 2008; Manski, 2013; Rodrik, 2015). This leads to a gap between formal econometric recommendations and empirical practice for over-identified models. [^2] That is, models which admit different estimators for the same parameter and consequently imply testable restrictions on the data distribution, c.f. Chen and Santos 2018.
Many textbook treatments of over-identified models (e.g. Hayashi, 2000; Stock and Watson, 2018; Davidson and Mackinnon, 2021) recommend researchers test whether the model’s over-identifying restrictions hold and use an efficient estimator when failing to reject. Experience reading economics papers and attending economics seminars, however, suggests that empirical practice often flouts these prescriptions. To document this, we examined empirical papers published in the American Economic Review between 2020 and 2024 which fit models using the generalized method of moments (GMM, Hansen, 1982) or related methods, obtaining a sample of 36 papers, of which 22 are over-identified. In this sample, we find that researchers commonly fit models using only a subset of the model’s implications, for instance fitting a fully parametric model by moment-matching rather than maximum likelihood. Researchers also commonly weight moments differently than would the optimal estimator (e.g. 14 of the over-identified papers in our sample report no efficiently-weighted specifications). Both practices yield inefficient estimators under correct model specification. Formal tests of over-identifying restrictions are notably scarce (only 3 papers in our sample report J-tests of overidentifying restrictions), though various “eyeball tests” are common, with researchers e.g. plotting the gap between data- and model-implied moments (25 papers in our sample).
@@
A similar perspective underlies a large and active recent literature in econometrics which studies the behavior and interpretation of popular estimators without assuming that the model is correct. This article provides a selective synthesis of this literature. While one can reasonably critique some of this literature as a reverse-engineering exercise, starting from estimators and working backwards in the hope of finding interpretable estimands (Mogstad and Torgovitsky, 2024), the perspective of this literature can also reasonably be summarized as “the purpose of an estimator is what it does.” [^3] Paraphrased from Beer (1985)—“The purpose of a system is what it does. There is, after all, no point in claiming that the purpose of a system is to do what is consistently fails to do.” That is, given that economic models are generally misspecified, it is important to understand how estimators actually behave in applications. Moreover, in our view, it can be reasonable to choose among estimators on this basis.
While reasonable, the common practice of choosing estimators based on misspecification concerns increases researcher degrees of freedom (Simmons et al., 2011). Different researchers estimating the same economic model on the same data may reach different conclusions due to differing misspecification concerns and subsequent estimation choices. Worse, dishonest researchers could use misspecification concern as a pretext for estimation choices which deliver a desired result, e.g. rejection of some null hypothesis.
As a first step towards quantifying the scope for such manipulation, and motivated by the wide range of weight-selection approaches observed in our survey of AER papers, we consider the choice of weight matrices for a given (over-identifying) set of moments. In models where the degree of misspecification is on the same order as sampling uncertainty (i.e. locally misspecified models) we prove that, asymptotically, the set of estimates obtainable at a given standard error is an interval centered at the efficient GMM estimate, with width proportional to the square root of Hansen (1982)’s J-statistic. Moreover, we show that a nefarious researcher can choose weights to engineer a t-statistic value of at least $\sqrt{J}$ for any null hypothesis. Thus, given a set of moments, J-statistics measure the scope for a nefarious researcher to “weight hack” their findings.
@@
    Researchers should use misspecification-robust standard errors, which excludes the usual GMM standard errors (Hansen, 1982) in over-identified models. Formulas are available for many problems, and when in doubt the bootstrap will often work.
@@
    Researchers estimating over-identified moment condition models should report J-statistics, since they summarize the range of results may be obtained via alternative weighting choices. J-statistics are thus important to report even if one dispenses with J-tests.
The following sections develop these points in detail, highlighting relevant references from the literature as we proceed. Section 2 defines what we mean by econometric and statistical models, illustrating with examples, and introduces corresponding notions of misspecification. Section 3 defines the estimands resulting from given estimators, and reviews some important estimand characterization results. Section 4 discusses misspecification-robust standard errors. Finally, Section 5 discusses the idea of local misspecification and derives what is to the best of our knowledge a new result, relating $J$-statistics to the range of estimates attainable under a standard error bound. Appendix A details a normal limit problem used to prove our $J$-statistic result, while Appendix B collects proofs for results stated in the main text.
@@ 2. What Do We Mean by “Misspecified?”
Suppose we observe a sample of $n$ observations $\{X_{1},...,X_{n}\}$ drawn from a distribution PP. [^4] Formally, we take $P$ to be a distribution over the infinite sequence $X^{\infty}=\{X_{1},X_{2},$...\} and suppose the sample is the first $n$ elements. Thus, our analysis covers both settings with iid data (in which case $P$ is a product measure) and cases with dependent data, subject to restrictions we note below. We are interested in using the data to learn an economic parameter $\theta\in\Theta,$ such as a counterfactual outcome or optimal policy. For the data to convey information about $\theta$, we must restrict which parameters can accompany a given data distribution. These restrictions could come from economic theory, from empirical evidence in other settings, or from ad-hoc restrictions chosen for analytical convenience. These restrictions imply an econometric model $\mathcal{M}\subseteq\Theta\times\Delta(\mathcal{X}),$ for $\Delta(\mathcal{X})$ the set of distributions on the sample space $\mathcal{X}.$ The model $\mathcal{M}$ collects pairs of $(\theta,P)$ that are consistent with our restrictions. $\mathcal{M}$ fully describes the relationship between the data distribution and the quantity of interest, and so determines the identification status of the parameter. [^5] While we discuss $\mathcal{M}$ as an econometric model given our focus on the economics literature, one might prefer another term, e.g. substantive model, since the same points arise when studying causal or counterfactual questions outside of economics. Spanos (2016) discusses related points under the heading of substantive adequacy. Specifically, the identified set for $\theta$ under data distribution $P$ is the set of parameter values consistent with $P$: $\Theta_{I}(P)=\left\{\theta\in\Theta:(\theta,P)\in\mathcal{M}\right\}.$ $\theta$ is point-identified if and only if $\Theta_{I}(P)$ is a singleton for all $P$ consistent with $\mathcal{M}$.
@@
Suppose we observe $X_{i}=(Y_{i},W_{i})$ drawn iid according to $P$. Here $Y_{i}\in\mathbb{R}$ is a scalar outcome while $W_{i}\in\mathcal{W}\subseteq\mathbb{R}$ is a continuous treatment. We believe that (i) each individual has potential outcome $Y_{i}(\cdot),$ where the observed outcome for individual $i$ is equal to that individual’s potential outcome evaluated at realized treatment, $Y_{i}=Y_{i}(W_{i})$; that (ii) treatment is randomly assigned with $W_{i}$ independent of $Y_{i}(\cdot)$; and that (iii) the potential outcomes are linear in $w,$
@@
This is a special case of the random coefficient models studied by e.g. Beran and Hall (1992). [^6] Note that if we take $W_{i}\in\{0,1\},$ rather than continuous, then this example is simply the potential outcomes model, $\alpha_{i}=Y_{i}(0)$ and $\beta_{i}=Y_{i}(1)-Y_{i}(0).$ Our parameter of interest is the average derivative $\theta=\mathbb{E}\left[\frac{\partial}{\partial w}Y_{i}(W_{i})\right],$ which describes the average effect of a marginal increase in $W_{i}.$ Under the model this average derivative can be written as $\mathbb{E}[\beta_{i}].$
Hence, under an additional regularity condition on a variance, which we shorthand with $\Xi=\operatorname{Var}(W_{i},\alpha_{i},\beta_{i})$, our econometric model is
$\mathcal{M}=\left\{\left({\mathbb{E}[\beta],\bigtimes_{i=1}^{\infty}\mathcal{L}(\alpha+\beta W,W)}\right):(\alpha,\beta)\indep W,\det(\Xi)\in(0,\infty)\right\}$
@@
$\mathcal{M}=\left\{(\vartheta(\psi),P):\psi\in\Psi,P\in\mathcal{P}^{R},\mathbb{E}_{P}[g(X_{i},\psi)]=0\right\}.$
@@
$\mathcal{P}=\left\{P\in\mathcal{P}^{R}:\text{ there exists }\psi\in\Psi\text{ such that }\mathbb{E}_{P}[g(X_{i},\psi)]=0\right\}$
Provided the identified set for $\psi$ is a singleton under each $P\in\mathcal{P}$—which in this context means that the moment equation $\mathbb{E}_{P}[g(X_{i},\psi)]=0$ is solved at a unique $\psi$—we can define $\psi^{*}:\mathcal{P}\to\Psi$ implicitly as the solution to the moment equations, $\mathbb{E}_{P}[g(X_{i},\psi)]=0,$ and can define the statistical parameter as $\theta^{*}(P)=\vartheta(\psi^{*}(P)).$ $\blacksquare$
@@
We say that the model is econometrically misspecified if the true $(\theta,P)$ pair falls outside the econometric model, $(\theta,P)\not\in\mathcal{M}.$ We assume the parameter $\theta$ remains well-defined under misspecification, which restricts the sorts of parameters we consider. In particular, our analysis is well-suited to e.g. counterfactual outcomes or causal effects, which may be defined (if not always estimated) non-parametrically. By contrast, for parameters which are ill-defined outside a specific model (e.g. preference parameters tied to a specific functional form), it makes little sense to contemplate a “true” $\theta$ outside of the model. Fortunately, in many economic analyses the objects of ultimate interest (e.g. counterfactuals, welfare, and optimal policies) are well-defined under much weaker assumptions than those often used to conduct estimation.
@@
Under this definition, the reason for misspecification $(\theta,P)\not\in\mathcal{M}$—incorrect relationship between $\psi$ and $P$ (i.e. the moments $g$), incorrect relationship between $\psi$ and $\theta$ (i.e. the function $\vartheta$), or some other source—is irrelevant. Indeed, since correct specification requires only that our model $\mathcal{M}$ allow the true $(\theta,P)$ pair, it is possible for the model to be correctly specified even if our underlying economic assumptions are incorrect and the model parameter $\psi$ does not have a well-defined “true” value. Similarly, the model is statistically misspecified at $P$ if there exists no value of $\psi$ that solves the moment conditions
@@
As noted by e.g. Breunig and Hoderlein (2018), the random coefficient model implies a large collection of over-identifying restrictions. For instance, the model implies that
@@
and thus that the conditional variance of $Y_{i}$ given $W_{i}$ must quadratic in $W_{i}.$ There are many ways for this (and other over-identifying restrictions) to fail, and thus for the model to be statistically misspecificed, $P\not\in\mathcal{P}$. [^7] The support of $W$ matters for this conclusion. If we instead consider binary $W_{i}\in\{0,1\}$ (and took e.g. $\theta=E[Y_{i}(1)-Y_{i}(0)],$ which is again equal to $E[\beta_{i}]$ under the model) then any $P$ with finite, nonzero $\operatorname{Var}_{P}(W_{i}Y_{i})$ and $\operatorname{Var}_{P}((1-W_{i})Y_{i})$ is consistent with $\mathcal{P}$.
@@
This article opts to focus on the latter approach.
@@ 2.1. An aside on the many meanings of models
Our definition of a statistical model follows a large literature in semiparametric statistics and econometrics, e.g. Chapter 25 of van der Vaart (2000) and Chen and Santos (2018). By contrast, phrases like “economic model” and “econometric model” are used in many different ways in the literature. Here we offer an incomplete survey and briefly relate some other definitions to ours. Our definition is in a sense minimal so long as one wishes to describe the properties of estimators for $\theta$. For instance, if we wish to characterize the set of parameter-value and bias pairs, $(\theta,\mathbb{E}_{P}\left[\hat{\theta}\right]-\theta),$ implied by any possible estimator $\hat{\theta}$ for $\theta,$ it is necessary and sufficient to know the set of possible values for the pair $(\theta,P).$ This, however, is precisely our definition of the econometric model.
A classical definition of a structural model is given by Koopmans and Reiersol (1950). They define a structure $Q$ as a particular joint distribution over latent and observed variables. [^9] Precisely, Koopmans and Reiersol (1950) define a structure as a distribution of latent variables and an equation linking observed and latent variables. A model $\mathcal{Q}$ in Koopmans and Reiersol (1950) is then a set of structures. The observed variables follow a distribution $P=P(Q)$, generated from a structure $Q\in\mathcal{Q}$. A structural parameter is a functional of the structure $\theta=\theta(Q)$. In our notation, a structural model in the Koopmans and Reiersol (1950) sense implies the econometric model
@@
Potential-outcomes notation and structural causal models also provide natural nesting models (Neyman, 1923/1990; Rubin, 1974; Pearl, 2009; Richardson and Robins, 2013; Heckman and Pinto, 2015). Many parameters of economic interest are functionals of the joint distribution $Q$ of all potential outcomes and treatments, suitably defined. [^10] In particular, structural equation models are often generative models for potential outcomes. Relating to Koopmans and Reiersol (1950), one could think of the joint distribution of potential outcomes and treatment variables as the structure. The distribution of potential outcomes then implies a distribution of the observed data $P(Q)$. Our approach nests this as well: $\mathcal{M}=\left\{{(\theta(Q),P(Q)):Q\in\mathcal{Q}}\right\}$ constitutes a model in our sense.
These notions are clearer and more versatile than some others. Some textbooks and reviews define an econometric model as deterministic economic relations plus stochastic disturbances (Intriligator, 1983; Woolridge, 2009), borrowing from the simultaneous equations literature. This usage is ambiguous, however, because the error term’s meaning is unclear: For instance, the equation
@@
may model linear, homogeneous causal effects $Y_{i}(w)=\alpha+\beta w+\epsilon_{i}$; or it may simply define $\beta$ as a statistical functional equal to the population OLS coefficient, leaving the joint distribution of $(Y,W)$ unrestricted. Under our terminology, the two cases correspond to distinct econometric models. [^11] The former describes a model from homogeneous and linear potential outcomes $\mathcal{M}=\left\{{(\beta,P):\bigtimes_{i=1}^{n}(Y_{i}(W),W_{i})\sim P,Y_{i}(w_{1})-Y_{i}(w_{2})=\beta(w_{1}-w_{2})}\right\}.$ The latter defines a model in which $\mathcal{M}=\left\{{\left({\frac{\operatorname{Cov}_{P}(W,Y)}{\operatorname{Var}_{P}(W)},P}\right):\bigtimes_{i=1}^{n}(Y_{i},W_{i})\sim P,\text{ $P$ has second moments}}\right\}.$

### v5

@@ Abstract.
In over-identified models, misspecification—the norm rather than exception—funda-
mentally changes what estimators estimate. Different estimators imply different estimands rather than different efficiency for the same target. A review of recent applications of generalized method of moments in the American Economic Review suggests widespread acceptance of this fact: There is little formal specification testing and widespread use of estimators that would be inefficient were the model correct, including the use of “hand-selected” moments and weighting matrices. Motivated by these observations, we review and synthesize recent results on estimation under model misspecification, providing guidelines for transparent and robust empirical research. We also provide a new theoretical result, showing that Hansen’s J-statistic measures, asymptotically, the range of estimates achievable at a given standard error. Given the widespread use of inefficient estimators and the resulting researcher degrees of freedom, we thus particularly recommend the broader reporting of J-statistics.
Keywords: Misspecification, Pseudo-true values, J-statistics
@@ 1. Introduction
Scientific and policy questions in economics frequently hinge on quantities such as counterfactual outcomes or effects on welfare. Recovering them often requires strong modeling assumptions. At the same time, many economists concede that economic models are at best approximately true (white1982maximum; hansen2008robustness; manski2013policy; rodrik2015econ). This leads to a gap between formal econometric recommendations and empirical practice for over-identified models. [^2] That is, models which admit different estimators for the same parameter and consequently imply testable restrictions on the data distribution, c.f. chen2018overidentification.
Many textbook treatments of over-identified models (e.g. hyashi2000econometrics; stock2018econometrics; davidson2003econometrics) recommend researchers test whether the model’s over-identifying restrictions hold and use an efficient estimator when failing to reject. Experience reading economics papers and attending economics seminars, however, suggests that empirical practice often flouts these prescriptions. To document this, we examined empirical papers published in the American Economic Review between 2020 and 2024 which fit models using the generalized method of moments (GMM, hansen1982gmm) or related methods, obtaining a sample of 36 papers, of which 22 are over-identified. In this sample, we find that researchers commonly fit models using only a subset of the model’s implications, for instance fitting a fully parametric model by moment-matching rather than maximum likelihood. Researchers also commonly weight moments differently than would the optimal estimator (e.g. 14 of the over-identified papers in our sample report no efficiently-weighted specifications). Both practices yield inefficient estimators under correct model specification. Formal tests of over-identifying restrictions are notably scarce (only 3 papers in our sample report J-tests or J-statistics of overidentifying restrictions), though various “eyeball tests” are common, with researchers e.g. plotting the gap between data- and model-implied moments (25 papers in our sample).
@@
A similar perspective underlies a large and active recent literature in econometrics which studies the behavior and interpretation of popular estimators without assuming that the model is correct. This article provides a selective synthesis of this literature. While one can reasonably critique some of this literature as a reverse-engineering exercise—starting from estimators and working backwards in the hope of finding interpretable estimands (mogstad2024IV)—the perspective of this literature can also reasonably be summarized as “the purpose of an estimator is what it does.” [^3] Paraphrased from beer1985diagnosing—“The purpose of a system is what it does. There is, after all, no point in claiming that the purpose of a system is to do what it consistently fails to do.” That is, given that economic models are generally misspecified, it is important to understand how estimators actually behave in applications. Moreover, in our view, it can be reasonable to choose among estimators on this basis.
While reasonable, the common practice of choosing estimators based on misspecification concerns increases researcher degrees of freedom (Simmonsetal2011). Different researchers estimating the same economic model on the same data may reach different conclusions due to differing misspecification concerns and subsequent estimation choices. Worse, dishonest researchers could use misspecification concern as a pretext for estimation choices which deliver a desired result, e.g. rejection of some null hypothesis.
As a first step towards quantifying the scope for such manipulation, and motivated by the wide range of weight-selection approaches observed in our survey of AER papers, we consider the choice of weight matrices for a given (over-identifying) set of moments. In models where the degree of misspecification is on the same order as sampling uncertainty (i.e. locally misspecified models) we prove that, asymptotically, the set of estimates obtainable at a given standard error is an interval centered at the efficient GMM estimate, with width proportional to the square root of hansen1982gmm’s J-statistic. Moreover, we show that a nefarious researcher can choose weights to engineer a t-statistic value of at least $\sqrt{J}$ for any null hypothesis. Thus, given a set of moments, J-statistics measure the scope for a nefarious researcher to “weight hack” their findings.
@@
    Researchers should use misspecification-robust standard errors, which excludes the usual GMM standard errors (hansen1982gmm) in over-identified models. Formulas are available for many problems, and when in doubt the bootstrap will often work.
@@
    Researchers estimating over-identified moment condition models should report J-statistics, since they summarize the range of results that may be obtained via alternative weighting choices. J-statistics are thus important to report even if one dispenses with J-tests.
The following sections develop these points in detail, highlighting relevant references from the literature as we proceed. Section 2 defines what we mean by econometric and statistical models, illustrating with examples, and introduces corresponding notions of misspecification. Section 3 defines the estimands resulting from given estimators, and reviews some important estimand characterization results. Section 4 discusses misspecification-robust standard errors. Finally, Section 5 reviews the idea of local misspecification and derives what is to the best of our knowledge a new result, relating J-statistics to the range of estimates attainable under a standard error bound. Appendix A details a normal limit problem used to prove our J-statistic result, while Appendix B collects proofs for results stated in the main text.
@@ 2. What Do We Mean by “Misspecified?”
Suppose we observe a sample of $n$ observations $X^{n}=\{X_{1},...,X_{n}\}$ drawn from a distribution PP. [^4] Formally, we take $P$ to be a distribution over the infinite sequence $X^{\infty}=\{X_{1},X_{2},$...\} and suppose the sample is the first $n$ elements. Thus, our analysis covers both settings with iid data (in which case $P$ is a product measure) and cases with dependent data, subject to restrictions we note below. We are interested in using the data to learn an economic parameter $\theta\in\Theta,$ such as a counterfactual outcome or optimal policy. For the data to convey information about $\theta$, we must restrict which parameters can accompany a given data distribution. These restrictions could come from economic theory, from empirical evidence in other settings, or from ad-hoc restrictions chosen for analytical convenience. These restrictions imply an econometric model $\mathcal{M}\subseteq\Theta\times\Delta(\mathcal{X}),$ for $\Delta(\mathcal{X})$ the set of distributions on the sample space $\mathcal{X}.$ The model $\mathcal{M}$ collects pairs of $(\theta,P)$ that are consistent with our restrictions. $\mathcal{M}$ fully describes the relationship between the data distribution and the quantity of interest, and so determines the identification status of the parameter. [^5] While we discuss $\mathcal{M}$ as an econometric model given our focus on the economics literature, one might prefer another term, e.g. substantive model, since the same points arise when studying causal or counterfactual questions outside of economics. spanos2016transforming discusses related points under the heading of substantive adequacy. Specifically, the identified set for $\theta$ under data distribution $P$ is the set of parameter values consistent with $P$: $\Theta_{I}(P)=\left\{\theta\in\Theta:(\theta,P)\in\mathcal{M}\right\}.$ $\theta$ is point-identified if and only if $\Theta_{I}(P)$ is a singleton for all $P$ consistent with $\mathcal{M}$.
@@
Suppose we observe $X_{i}=(Y_{i},W_{i})$ drawn iid. Here $Y_{i}\in\mathbb{R}$ is a scalar outcome while $W_{i}\in\mathcal{W}\subseteq\mathbb{R}$ is a continuous treatment. We believe that (i) each individual has potential outcome $Y_{i}(\cdot),$ where the observed outcome for individual $i$ is equal to that individual’s potential outcome evaluated at realized treatment, $Y_{i}=Y_{i}(W_{i})$; that (ii) treatment is randomly assigned with $W_{i}$ independent of $Y_{i}(\cdot)$; and that (iii) the potential outcomes are linear in $w,$
@@
This is a special case of the random coefficient models studied by e.g. beran1992estimating. [^6] Note that if we take $W_{i}\in\{0,1\},$ rather than continuous, then this example is simply the potential outcomes model, $\alpha_{i}=Y_{i}(0)$ and $\beta_{i}=Y_{i}(1)-Y_{i}(0).$ Our parameter of interest is the average derivative $\theta=E\left[\frac{\partial}{\partial w}Y_{i}(W_{i})\right],$ which describes the average effect of a marginal increase in $W_{i}.$ Under the model this average derivative can be written as $E[\beta_{i}].$
Under an additional regularity condition on a variance, which we shorthand with $\Xi=\operatorname{Var}(W_{i},\alpha_{i},\beta_{i})$, our econometric model is
$\mathcal{M}=\left\{\left({E[\beta],\bigtimes_{i=1}^{\infty}\mathcal{L}(\alpha+\beta W,W)}\right):(\alpha,\beta)\indep W,\det(\Xi)\in(0,\infty)\right\}$
@@
$\mathcal{M}=\left\{(\vartheta(\psi),P):\psi\in\Psi,P\in\mathcal{P}^{R},E_{P}[g(X_{i},\psi)]=0\right\}.$
@@
$\mathcal{P}=\left\{P\in\mathcal{P}^{R}:\text{ there exists }\psi\in\Psi\text{ such that }E_{P}[g(X_{i},\psi)]=0\right\}.$
Provided the identified set for $\psi$ is a singleton under each $P\in\mathcal{P}$—which in this context means that the moment equation $E_{P}[g(X_{i},\psi)]=0$ is solved at a unique $\psi$—we can define $\psi^{*}:\mathcal{P}\to\Psi$ implicitly as the solution to the moment equations, $E_{P}[g(X_{i},\psi)]=0,$ and can define the statistical parameter as $\theta^{*}(P)=\vartheta(\psi^{*}(P)).$ $\blacksquare$
@@
We say that the model is econometrically misspecified if the true $(\theta,P)$ pair falls outside the econometric model, $(\theta,P)\not\in\mathcal{M}.$ We assume the parameter $\theta$ remains well-defined under misspecification, which restricts the sorts of parameters we consider. In particular, our analysis is well-suited to, e.g., average counterfactual outcomes or causal effects, which may be defined (if not always estimated) non-parametrically. By contrast, for parameters which are ill-defined outside a specific model (e.g., preference parameters tied to a specific functional form), it makes little sense to contemplate a “true” $\theta$ outside of the model. Fortunately, in many economic analyses the objects of ultimate interest (e.g., counterfactuals, welfare, and optimal policies) are well-defined under much weaker assumptions than those often used to conduct estimation.
@@
For instance, it might be that the moment conditions are not solved by any parameter value $\psi$ consistent, through $\vartheta$, with the true counterfactual outcome $\theta$. Under this definition, the reason for misspecification $(\theta,P)\not\in\mathcal{M}$—incorrect relationship between $\psi$ and $P$ (i.e. the moments $g$), incorrect relationship between $\psi$ and $\theta$ (i.e. the function $\vartheta$), or some other source—is irrelevant. Indeed, since correct specification requires only that our model $\mathcal{M}$ allow the true $(\theta,P)$ pair, it is possible for the model to be correctly specified even if our underlying economic assumptions are incorrect and the model parameter $\psi$ does not have a well-defined “true” value. Similarly, the model is statistically misspecified at $P$ if there exists no value of $\psi$ that solves the moment conditions
@@
As noted by e.g. breunig2018specification, the random coefficient model implies a large collection of over-identifying restrictions. For instance, the model implies that
@@
and thus that the conditional variance of $Y_{i}$ given $W_{i}$ must be quadratic in $W_{i}.$ There are many ways for this (and other over-identifying restrictions) to fail, and thus for the model to be statistically misspecificed, $P\not\in\mathcal{P}$. [^7] The support of $W_{i}$ matters for this conclusion. If we instead consider binary $W_{i}\in\{0,1\}$ (and took e.g. $\theta=E[Y_{i}(1)-Y_{i}(0)],$ which is again equal to $E[\beta_{i}]$ under the model) then any $P$ with finite, nonzero $\operatorname{Var}_{P}(W_{i}Y_{i})$ and $\operatorname{Var}_{P}((1-W_{i})Y_{i})$ is consistent with $\mathcal{P}$.
@@
This article focuses on the latter approach.
@@ 2.1. An aside on the many meanings of models
Our definition of a statistical model follows a large literature in semiparametric statistics and econometrics, e.g. Chapter 25 of van2000asymptotic and chen2018overidentification. By contrast, phrases like “economic model” and “econometric model” are used in many different ways in the literature. Here we offer an incomplete survey and briefly relate some other definitions to ours. Our definition is in a sense minimal so long as one wishes to describe the properties of estimators for $\theta$. For instance, if we wish to characterize the set of parameter-value and bias pairs, $(\theta,E_{P}[\hat{\theta}]-\theta),$ implied by any possible estimator $\hat{\theta}$ for $\theta,$ it is necessary and sufficient to know the set of possible values for the pair $(\theta,P).$ This, however, is precisely our definition of the econometric model.
A classical definition of a structural model is given by koopmans1950identification. They define a structure $Q$ as a particular joint distribution over latent and observed variables. [^9] Precisely, koopmans1950identification define a structure as a distribution of latent variables and an equation linking observed and latent variables. A model $\mathcal{Q}$ in koopmans1950identification is then a set of structures. The observed variables follow a distribution $P=P(Q)$, generated from a structure $Q\in\mathcal{Q}$. A structural parameter is a functional of the structure $\theta=\theta(Q)$. In our notation, a structural model in the koopmans1950identification sense implies the econometric model
@@
Potential outcomes notation and structural causal models also provide natural nesting models (splawa1990application; rubin1974estimating; pearl2009causality; richardson2013single; heckman2015causal). Many parameters of economic interest are functionals of the joint distribution $Q$ of all potential outcomes and treatments, suitably defined. [^10] In particular, structural equation models are often generative models for potential outcomes. Relating to koopmans1950identification, one could think of the joint distribution of potential outcomes and treatment variables as the structure. The distribution of potential outcomes then implies a distribution of the observed data $P(Q)$. Our approach nests this as well: $\mathcal{M}=\left\{{(\theta(Q),P(Q)):Q\in\mathcal{Q}}\right\}$ constitutes a model in our sense.
These notions are clearer and more versatile than some others. Some textbooks and reviews define an econometric model as deterministic economic relations plus stochastic disturbances (intriligator1983economic; woolridge2009introductory), borrowing from the simultaneous equations literature. This usage is ambiguous, however, because the error term’s meaning is unclear: For instance, the equation
@@
may model linear, homogeneous causal effects $Y_{i}(w)=\alpha+\beta w+\epsilon_{i}$; or it may simply define $\beta$ as a statistical parameter equal to the population OLS coefficient, leaving the joint distribution of $(Y,W)$ unrestricted. [^11] Foundational works in econometrics such as haavelmo1943simultenous; haavelmo1944econometrics and koopmans1950identification clearly distinguish these possibilities. See e.g. pearl2015haavelmo for discussion. Under our terminology, the two cases correspond to distinct econometric models. [^12] The former describes a model of homogeneous and linear potential outcomes $\mathcal{M}=\left\{{(\beta,P):\bigtimes_{i=1}^{n}(Y_{i}(W),W_{i})\sim P,Y_{i}(w_{1})-Y_{i}(w_{2})=\beta(w_{1}-w_{2})}\right\}.$ The latter defines a model in which $\mathcal{M}=\left\{{\left({\frac{\operatorname{Cov}_{P}(W,Y)}{\operatorname{Var}_{P}(W)},P}\right):\bigtimes_{i=1}^{n}(Y_{i},W_{i})\sim P,\text{ $P$ has second moments}}\right\}.$